Please help me please. Solve the inequality analytically.
x square - 2x - 35 < 0
Thank you

Answer by J2R2R(94)

x^2 - 2x - 35 = (x-7)(x+5)

For this to be less than zero, think of the graph which crosses the x-axis at -5 and 7 (where x^2 - 2x - 35 = 0) and between these two points the graph is below the x-axis i.e. negative in value.

Therefore the range for x^2 - 2x - 35 < 0 is -5 < x < 7